OpenStudy (anonymous):

I can't figure out how to set up this inequality problem. Mario makes 6 free throw shots out of his first 10 attempts. What is the least number of consecutive shots Mario must make to increase his percentage from 60% to at least 70%? ~ 1 shot ~ 4 shots ~ 10 shots ~ 14 shots

OpenStudy (anonymous):

I'm pretty sure that one of the fractions will be 6/10 since he made 6 shot out of his first 10 attempts.

OpenStudy (mitu12):

currently: 6/10 = 60% letting x = additional free throws to make 70%: (6 + x) / (10 + x) = 70% solving for x: (6 + x) / (10 + x) = 7/10 cross multiply: 60 + 10x = 70 + 7x 3x = 10 x = 3 1/3 since he can't make 1/3 of a shot, the least number of consecutive shots to get to 70% would be 4.

OpenStudy (anonymous):

I see now. Thank you very much. :)

OpenStudy (mitu12):

no problem